For the right triangle shown, the lengths of two sides are given. Find the third side. Leave your answer in simplified, radical form.


a = 4, b = 4, c =???

Answer by Boreal(15235)

a^2+b^2=c^2
4^2+4^2=c^2=32
c=sqrt(32)=sqrt(16)*sqrt(2)=4 sqrt (2)